A baseball player strikes a ball a 0.145kg ball which reaches a height of 115m, determine the velocity it left the bat at if contact occured at 1.5m off the ground.


1
Expert's answer
2022-04-27T13:34:11-0400

"v=\sqrt{2g(H-h)}=47.2~\frac ms."